INAF LBC+MODS+LUCI queue observing, 2012 October 19-20

Observer: M. Faccini, E. Sani, K. Boutsia
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: M. Edwars
Instrument: LUCI1

Summary

Clouds in the horizon and thin scattered cirrus all over the sky. Not photometric at the beginning of the night.
We start with a Hunt_8G LUCI1 project before we run the transit project
We had a lot of problems with the transit project - main one was that the finding chart was missing and we could not identify the target from the chart in the proposal. Reference stars were outside of the field. After 1hr of effort, we abort and we move to the next targets - Anyway, weather was partly cloudy - at 5:30UT we had to close the dome because of clouds.
The Turatto target is too low (>2.0 airmass) and the offsets are too large, so it looses the guide star. We abort.
Project Hunt_8A has a faint target, impossible to be recognised without a blind offset. We abort.
La Franca Projects A1 and A2 were executed and completed

At the end of night we obtained luci flats + arcs for Cresci and LaFranca. Dark frames for this data can be found in UT20121221.

ATTENTION: All LUCI data from tonight do not have reliable filter+grating information in the header. Check log to see what filter was used for each block of data

Data Summary

Project_RUN Objects Filter ExpTime Notes
Hunt_8G GRB080325 HKspec, 200H+K 20min we are not sure that the target is within the slit
Sozzeti WASP33 HKspec, 200H+K - we are not able to do the mask acquisition because the field does not seem to match the mask - after 1hr of effort we abort
Turatto_07A SN2009ip HKspec - target too low and we loose guide star in one offset - abort
Cresci_28A DSF2237-A H, 210_zJHK 30m flat +arc +Tell
La Franca 2MSXJ06+32 zJspec, 200 H+K 1hr flat + arc +Tell - COMPLETED
Hunt_8A GRB020127 HKspec, 200H+K - target too faint to identify in the acq image - we need a blind offset -abort
La Franca 2MSXJ09+45 zJspec, 200H+K 1hr flat + arc +Tell - COMPLETED

Details

01:39 Start Hunt Program G

1:38 executeLUCIScipt.sh GRB080325acquisition.script. luci.20121020.0003.fits luci.20121020.0004.fits
  • Target is barely visible and we are not sure it is the correct one.
  • we try with the blind offset

1:51 executeLUCIScipt.sh GRB080325acquisition_blind.script
  • again we do not see the target in the slit... it should show - even though it would be very very faint.

2:09 executeLUCIScipt.sh GRB080325science_HKspec.script luci.20121020.0010.fits luci.20121020.0011.fits
  • there is no visible spectrum ... we want to take the slit off and see if the target is where it is supposed to be.

2:24 the filter wheel FAILED - the transit targets starts to be visible - we abort Hunt project

2:34 the T.O. is slewing to the target because we cannot use LUCI

2:44 preset to Sozzetti 22
  • NO filters in the script
  • there is no Finding Chart and we are not sure what we are looking at
  • using the chart annex to the proposal we try to figure wich is the target star and we start looking for the reference stars
  • the MOS acq software gives us the first star outside the field ... obviously not correct... since we have reference only in the proposal, and no useful finding chart, we can not look for them.
  • variable seeing, clouds
  • imgs taken luci.20121020.0012.fits luci.20121020.0028.fits

3:42 we invested 1hr in this project and since we cannot find the targets in a reliable way, the beginning time for the project was already passed, the weather was partly cloudy, we decide to switch targets

03:45 slew to Turatto 7A / 7B
  • we do not see acq script for the spectra and we start with the imaging SN2009ip_oct2012_H_ImagingSc luci.20121020.0031.fits luci.20121020.0042.fits
  • the guide star in the imaging scripts is different than in the acq script - corrected
  • the offsets in the imaging scripts are too large and the guide star goes out of range.
  • we modify acq imaging script for spectroscopy acq and we decide to start with the spectra

4:05 Long slit acquisition - for SN2009qi - Field of view hardly recognisable, probably due to clouds

4:18 we start specHK science script for Turatto7A SN2009ip_HK.spec luci.20121020.0048.fits luci.20121020.0050.fits
  • offsets along the slit are too large and the guide star is out of reach.

04:25 abort Turatto 7A / 7B

4:36 start Cresci 28A
  • aborted, wrong camera - during this run both acq and science have to be done with the N1.8camera

4:43 Cresci start again after correcting mask ID and camera N1.8 - wrong ID number

5:17 Start Cresci science script sci_campoA_H_1.txt luci.20121020.0057.fits
  • we are loosing the guide star because of clouds

5:27 Closing Dome

6:00 open Dome again for observations - restart Cresci A acquisition

6:21 we start specs. Campo A science H sci_campoA_H_1.txt luci.20121020.0062.fits luci.20121020.0067.fits

6:59 slew to the Telluric to obtain right air mass tell_int_campoA_H.txt luci.20121020.0070.fits luci.20121020.0077.fits

7.19 slew to La Franca A1 OB_LaFranca_A1_acq.txt luci.20121020.0078.fits luci.20121020.0080.fits

7:24 repeat preset after unwrap. collimation is taking long because of clouds

7:28 repeat preset due to collimation

7:30 repeat preset

7:41 problems setting the zJspec through the slit. it has to be set by hand by the support astronomer

7:49 repeat preset

7:50 executeLUCIScipt.sh OB_LaFranca_A1_spe.txt luci.20121020.0084.fits luci.20121020.0091.fits

8:40 Slew to STD for LaFranca

8:55 Science frames for STD
  • the exposure time is 15sec and the counts are high OB_LaFranca_STD_spe.txt luci.20121020.0097.fits
  • we lower the exptime to 5sec and we repeat OB_LaFranca_STD_speSHORT.txt luci.20121020.0099.fits luci.20121020.0102.fits
  • LaFranca_A1 completed

9:05 Slew to Hunt_8A
  • we execute acquisition script - the target should be visible GRB020127acquisition.script
  • it is not visible so we increase exposure time to 15 x 10
  • again not visible - we increase exposure time to 20 x 10
  • again not visible - we change filter from H to Ks because on the finding chart gives Ksmag=18.5 ( on proposal was Ksmag=20)
  • for exptime 5 x 5 not visible - we increase exptime to 15x 10 - not visible again
  • we cannot center the slit on this object - we need a reference star in the slit - or a blind offset luci.20121020.0103.fits luci.20121020.0112.fits

10:05 Hunt_8A aborted

10:08 We slew to LaFranca A2

10:25 science OB for LaFranca A2 OB_LaFranca_A2_spe.txt luci.20121020.0118.fits luci.20121020.0125.fits

11:29 End of science for tonight - we slew to the LaFranca STD to practice longslit acquisition with LUCI OB_LaFranca_STD_speSHORT.txt luci.20121020.0131.fits luci.20121020.0132.fits

  • we repeated the STD observations and also to source A2 LS acquisition OB_LaFranca_A2_spe.txt luci.20121020.0139.fits luci.20121020.0144.fits
  • after that we obtain flats and arcs for Cresci and LaFranca

-- RobertoSpeziali - 20 Oct 2012
Topic revision: r12 - 05 Dec 2012, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback